# Generative Verifiers: Reward Modeling as Next-Token Prediction

### v1

@@ 3.3 Chain-of-Thought Verifiers (GenRM-CoT)
Instead, we use reference-guided grading (Zheng et al., 2024) to improve the quality of synthetic rationales. Specifically, we provide a reference solution in addition to the problem and solution to verify (see Table A.2), making it easier for an LLM to point out any reasoning error in the provided solution. Here, a reference solution is defined as any model-generated solution that manages to arrive at the correct final answer. Note that reference-guided grading can only be used during training, as we do not have reference solutions for test problems.
@@
Figure 4: An example where GenRM-CoT catches a subtle mistake that the discriminative verifier is unable to catch. The candidate solution did not convert 90 minutes into 1.5 hours before dividing it by 7.5. However, the discriminative verifier was not able to detect this mistake likely because the solution does still appear to produce a valid-sounding percentage 90/7.5 = 12. Our proposed GenRM-CoT model is able to identify this mistake using step-by-step generative verification. The full verification output can be found in Table C.11.
4 Experiments
The goal of our experiments is to demonstrate the efficacy of next-token prediction compared to other approaches for training verifiers. To this end, we compare GenRM and standard verifiers on a number of reasoning problems with the goal of answering the following research questions:
  • How does GenRM compare to standard discriminative verifiers in reasoning domains?
  • Does unified training of GenRM improve generation and verification performance?
  • Can GenRM effectively utilize CoT reasoning and test-time compute to improve its performance?
  • How does GenRM scale with data and model size?
Problem: Digimon had its 20th anniversary. When it came out John was twice as old as Jim. If John is 28 how old is Jim?
We know that 2x - 20 = 28 (since John’s current age is 28)
2x = 48
x = 24
So, when Digimon came out, Jim was 24. Since it’s been 20 years, Jim is now 24 + 20 = 44 years old. The answer is 44.
Discriminative RM score: 0.9778
GenRM-CoT: Let’s verify step by step.
$\cdots$
Step 2: 20 years ago, John’s age was 2x - 20 (current age - 20 years) and Jim’s age was x - 20 (current age - 20 years). (Correct)
$\cdots$
Step 6: So, when Digimon came out, Jim was 24. Since it’s been 20 years, Jim is now 24 + 20 = 44 years old. (Incorrect)
**Expected Answer:**
$\cdots$
Verification: Is the solution correct (Yes/No)? No
GenRM-CoT (Majority Voting) score: 0.0000
Figure 5: What happens when GenRM does not catch the mistake in its first pass? At the end of the step-by-step verification, GenRM-CoT tends to output a correct/incorrect token summarizing if the final answer seems wrong, similar to a direct GenRM that only predicts Yes/No. If GenRM deems the final answer incorrect, it will attempt to solve the problem again using a reasoning path different from the solution. In the example above, even though the step-by-step verification fails to find any mistake, GenRM gives the problem another try and correctly verifies the solution. Full verification output in Figure C.1. This behavior emerges from training on synthetic rationales generated using the prompting format shown in Table A.2.
Figure 6: Comparing generative verifiers with other approaches. GenRM generally outperforms other methods, and more closely matches the Best-of-N performance with oracle verifier ( Pass@N). Moreover, GenRM with majority voting is able to effectively utilize chain-of-thought reasoning and test-time compute to further improve over direct GenRM verifier.
Tasks and Data Generation. We focus on the following reasoning tasks (more details in Appendix A):
  • Last Letter Concatenation (Wei et al., 2022): Given a list of words, the task is to concatenate the last letters of each word (for instance, "Noah Paul Elisha Rebecca" $\rightarrow$ "hlaa"). We train verifiers on lists of lengths $2-4$, and evaluate the verifier on the out-of-distribution (OOD) setting of length 6.
  • Word Sorting (Suzgun et al., 2022) Given a list of words, sort them in alphabetical order. For a sorting task, verification is clearly easier than generating the correct solution. Similar to the last letter concatenation task, we train verifiers on up to 4-words, and evaluate length-generalization performance on 5 word examples.
  • GSM8K (Cobbe et al., 2021) is a widely-used dataset to evaluate grade-school math reasoning capabilities of LLMs. For training, we use at most 16 correct and 16 incorrect solutions per problem. We evaluate the verifier performance on 16 solutions per problem in the test set.
Baselines. We compare GenRM to the following standard approaches for verification:
  • Discriminative (standard) RM (ORM) (Cobbe et al., 2021): The prevalent approach for training verifiers for test-time re-ranking on reasoning tasks, as discussed in §2, serves as our main baseline.
  • LLM-as-a-Judge (Zheng et al., 2024): This approach uses an off-the-shelf pretrained LLM for verification. To do so, we use a CoT prompt to produce 32 verification rationales that is used for correctness prediction and pick the majority vote correctness answer.
Evaluation protocol. Following prior work (Cobbe et al., 2021; Lightman et al., 2023), we primarily use Best-of-N performance (in terms of the percentage of problems solved) of a fixed LLM (generator) (§2) with learned verifiers, and report average performance on the test set. Best-of-N evaluates what fraction of solutions chosen by the verifier are correct. For some results, we also report RM accuracy on the test set or a held-out validation set, which measures whether the verifier accurately classifies incorrect and correct solutions. While these two metrics are often correlated, RM accuracy only evaluates the verifier’s point-wise accuracy, while Best-of-N evaluates its group-wise ranking performance.
Models. For training discriminative RMs and GenRM, we use open-weights Gemma models (Team et al., 2024), specifically Gemma-2B for algorithmic tasks, and Gemma 2B, 7B, and 9B for GSM8K. For solution generation as well as LLM-as-a-Judge, we use Gemma 2B for algorithmic tasks and Gemini 1.0 Pro (Team et al., 2023) for GSM8K.
Hyperparameters. By default, all GenRM experiments use unified training for verification with solution generation (5), with $\lambda=1/3$ for algorithmic tasks and $\lambda=1/4$ for GSM8K. We use the label ‘Verification Only’ to indicate GenRMor GenRM-CoT verifiers trained using only verification data ($\lambda=0$). See Appendix Appendix B for more details.
Verification CoT Rationale Generation. For training rationales, we generated ground-truth rationales for word sorting and last letter concatenation algorithmically, as shown in Table A.1. On GSM8K, we generated rationales using Gemini 1.0 Pro with reference-guided grading (Zheng et al., 2024), with the prompt outlined in Table A.2.
Figure 7: SFT on correct solutions enhances verification performance, both for GenRM that directly predicts Yes/No, and GenRM-CoT that uses verification CoT rationales, across all tasks that we consider. Here, ‘Verification Only’ corresponds to training GenRM or GenRM-CoT only on verification data, by setting $\lambda=0$ in (5).
Figure 8: Unifying generation and verification boosts GenRM’s generation performance. Compared to SFT only on correct solutions, jointly training on generation data and CoT verification data results in improved Best-of-N performance (with oracle verifier) for the generator. Furthermore, we observe that the generation performance of GenRM-CoT exhibits more improvement on algorithmic tasks than on GSM8K, likely because those tasks use ground truth verification rationales while GSM relies on synthetic LLM-generated rationales, which may be inaccurate.
4.1 Comparing GenRM with Prior Verification Approaches
GenRM, which directly predicts Yes/No token for verification, can match or outperform the discriminative RM and other approaches on all the three tasks, as shown in Figure 6. This shows that the next-token-prediction loss allows GenRM to tap into the capabilities of pretrained Gemma models more effectively.
CoT Reasoning Improves Verification. GenRM-CoT, which combines chain-of-thought with majority voting, further improves the performance over GenRM.
In particular, on the two algorithmic tasks with oracle verification CoTs, GenRM-CoT closely matches the oracle verifier performance (Pass@N). On GSM8K, GenRM-CoT consistently outperforms all other methods, even though the training CoT rationales (generated with Gemini 1.0 Pro) may contain errors. Qualitatively, GenRM-CoT is able to detect subtle reasoning errors that are missed by discriminative verifiers (see Figure 2, 5,4).
4.2 Unifying Generation and Verification
Unifying solution generation with verification, as done by GenRM using the next-token-prediction objective, consistently improves verification performance across all tasks, as illustrated in Figure 7. This improvement is observed for both direct and CoT-based generative verifiers, suggesting that teaching the verifier to imitate correct solutions generally helps.
Notably, incorporating CoT verification data into the generator’s training mix leads to better solution generation performance for the GenRM-CoT verifier itself, as evidenced in Figure 8 by the improved Best-of-N scores with the oracle verifier (Pass@N). This suggests that teaching a generator to perform verification based on next-token prediction can deepen its understanding of the generation process itself. Overall, the above results indicate the unifying solution generation and verification is mutually beneficial.
Figure 9: Scaling Inference-time Compute for Verification. By posing reward modeling as next-token prediction, GenRM can utilize the powerful toolbox from generative LLMs, such as Chain-of-Thought and Majority Voting, to turn additional test-time compute into higher Best-of-N accuracy. The more capable the base model, the more effective inference-time compute scaling seems to become. Here, the horizontal line corresponds to performance of GenRM-CoT verifier with greedy decoding.
Figure 10: Model Scaling for Generative Verifiers. On GSM8K, we scale model capacity using Gemma (Team et al., 2024a) 2B. 7B, 9B, and observe positive scaling trends both for GenRM (direct) and GenRM-CoT, both for Best-of-N performance as well as reward modeling accuracy on the test set.
Figure 11: Data scaling for GenRM-CoT on GSM8K. We observe that both the RM accuracy and Best-of-N performance improve as we scale up the number of rationales per solution and solutions per problem. When adding more solutions, we use 4 rationales per solution. Here, we compute GenRM-CoT scores with CoT rationales generated using greedy decoding, as discussed in (6).
Figure 12: Data efficiency for GenRM vs a standard verifier. Overall we observe a positive but moderate scaling trend for all methods, with GenRM consistently outperforming the discriminative verifier in all data regimes. In addition, we find that adding a solution generation loss ($\lambda>0$) can further help GenRM, with $\lambda=1/4$ being a good value for GSM.
Figure 13: Using reference guidance in the prompt for synthetic rationale generation is crucial for GenRM-CoT to perform well.
4.3 Scaling Data, Model Size, and Inference-time Compute
Scaling Test-Time Compute with GenRM-CoT can be done by sampling multiple CoTs and conduct majority voting, as described in (7). As shown in Figure 9, GenRM-CoT verifier’s performance scales gracefully with greater number of votes at test time, under all three Gemma scales (2B, 7B, 9B), outperforming greedy decoding performance within 4 votes. Across model scales, the finetuned GenRM-CoT verifier outperforms LLM-as-a-Judge, which also utilizes the same CoT approach and number of majority votes, but prompts a more capable Gemini 1.0 Pro model than Gemma models.
Scaling model size. In Figure 10, we show that the performance of generative verifiers scales up positively with an increase in model capacity. The model scaling experiments use the fixed dataset (32 solutions per problem, and for CoT GenRM, 4 rationales per solution). The results show that bigger models are able to learn more from the same data, which matches what we expect from scaling model parameter counts under the next-token prediction loss.
Data scaling for CoT verifiers. GenRM-CoT allows for an additional axis of data scaling, which is absent in standard verifiers: scaling the number of rationales per solution. On GSM8K, we find that using multiple rationales per solution has a substantial effect on the performance of generative verifiers. We suspect that this is because model-generated, synthetic rationales are noisy in this case, such that training on multiple rationales per solution and the associated “ensembling” effect prevents the training procedure from overfitting to this noise and spurious correlations. See Figure 11 for detailed comparisons. Both RM Accuracy and Best-of-N Accuracy scales positively on both data axis, with the number of rationales per solution having a bigger effect.
Data scaling for direct GenRM. GenRM trained on verification only data still outperforms standard verifiers as we increase the number of solutions per problem on GSM8K (Figure 13, left), demonstrating the effectiveness of casting verification as a next-token prediction problem. Across all data scales, unified training with solution generation data further boosts the performance of GenRM verifiers, as already discussed in §4.2. Moreover, the optimal loss coefficient for solution generation data in GenRM follows an inverted U-shape: adding too little or too much negatively impacts verification, while intermediate values yield the best results (Figure 13, right).
4.4 Impact of Synthetic Rationale Quality
Our results on GSM8K results indicate that GenRM-CoT verifier can outperform discriminative and direct GenRM verifiers even without human-written rationales, highlighting the potential of LLM-generated rationales. However, the quality of these synthetic rationales does matter, as shown in Figure 13. Using reference-guided grading during rationale generation significantly improves performance (91.7% with guidance vs. 87.8% without for Gemma-7B verifiers), indicating that LLMs are better at identifying reasoning errors when they have a reference solution for comparison. Importantly, achieving our result does not require a more capable model for generating verification rationales: we use the same model (Gemini 1.0 Pro) to generate both solutions to verify and synthetic rationales in the training data.

### v2

@@ 3.3 Chain-of-Thought Verifiers (GenRM-CoT)
Instead, we use reference-guided grading (Zheng et al., 2024) to improve the quality of synthetic rationales, which we filter using their verification correctness. Specifically, we provide a reference solution in addition to the problem and solution to verify (see Table A.2), making it easier for an LLM to point out any reasoning error in the provided solution. Here, a reference solution could be any model-generated solution that arrives at the correct final answer. Note that reference-guided grading can only be used during training, as we do not have reference solutions for test problems.
Problem: A line is parameterized by $\begin{pmatrix}x\\ y\end{pmatrix}=\begin{pmatrix}8\\ -1\end{pmatrix}+t\begin{pmatrix}2\\ 3\end{pmatrix}$. The equation of the line can be expressed in the form $y=mx+b.$ Enter the ordered pair $(m,b).$
Solution: We have $x=8+2t$ and $y=-1+3t$. Solving the first equation for $t$, we have $t=\frac{x-8}{2}$. Substituting this into the second equation, we get $y=-1+3\left(\frac{x-8}{2}\right){\color[rgb]{1,0,0}=\frac{3}{2}x-\frac{23}{2}}$. Thus, $m=\frac{3}{2}$ and $b=-\frac{23}{2}$. So, the answer is $\boxed{\left(\frac{3}{2},-\frac{23}{2}\right)}$.
Discriminative RM correctness score: 0.827
GenRM-CoT: Let’s verify step by step. $\cdots$
**Step 3:**
Solution: Substituting this into the second equation, we get $y=-1+3\left(\frac{x-8}{2}\right)=\frac{3}{2}x-\frac{23}{2}$.
Expected Answer: Substituting this into the second equation, we get $y=-1+3*(x-8)/2={\color[rgb]{0,0,1}-1+(3x-24)/2=-1+(3x/2)-12=(3x/2)-13}$.
**Verification:** No
$\cdots$
Verification: Is the answer correct (Yes/No)? No
GenRM-CoT (Majority Voting) score: 0.438
Figure 4: An example on MATH where GenRM-CoT (trained only on GSM) detects a reasoning error. The solution made a mistake in simplifying an intermediate step. Both Discriminative RM and GenRM-CoT models have only been trained on GSM8K. In this case, discriminative RM fails to classify the solution as incorrect, whereas GenRM-CoT utilizes chain of thoughts to catch this mistake. See Table E.14 for details.
4 Experiments
In this section, we evaluate the efficacy of next-token prediction and chain-of-thought reasoning for verification compared to standard verification approaches. To this end, we compare GenRM and standard verifiers on a number of reasoning tasks to answer the following questions: (1) How does GenRM compare to discriminative verifiers and other approaches? (2) Does unified training of GenRM improve generation and verification performance? (3) Can GenRM effectively utilize CoT reasoning to improve its performance? (4) How does GenRM scale with model size and inference-time compute?
Tasks. We focus on the following tasks and put details about data generation in Appendix A:
  • Algorithmic reasoning. We use two difficult string manipulation tasks, namely Last Letter Concatenation (Wei et al., 2022) and Word Sorting from Big-Bench (Suzgun et al., 2022). We train verifiers on word lists of length {2,3,4}, and evaluate their generalization on length {5,6}.
  • Math reasoning. We train grade-school math verifiers on the GSM8K dataset from Cobbe et al. (2021) that popularized test-time verification. We evaluate these verifiers on the GSM8K test set as well as their easy-to-hard generalization on much harder MATH dataset (Hendrycks et al., 2021), using the same held-out set of 500 MATH problems as Lightman et al. (2023).
Baselines. We compare GenRM to the following verification approaches:
  • Discriminative RM (Cobbe et al., 2021) or ORM is the prevalent approach for training verifiers for test-time re-ranking on reasoning tasks (§2), and serves as our main baseline.
  • LLM-as-a-Judge (Zheng et al., 2024) uses an off-the-shelf pretrained LLM for verification. To do so, we use a CoT prompt to produce 32 verification rationales that is used for correctness prediction and pick the majority-vote correctness answer.
  • DPO (Rafailov et al., 2024): Following Hosseini et al. (2024), we use this preference optimization approach for training verifiers on preference pairs with incorrect and correct solutions.
Note that self-consistency and test-time verification are complementary approaches, and can be often combined via weighted self-consistency to further boost performance, as shown in Figure 5.
Evaluation protocol. Following Cobbe et al. (2021); Lightman et al. (2023), we primarily use Best-of-N performance in terms of the percentage of problems solved using a fixed generator (§2) with learned verifiers, and report average accuracy on the test set. We also report test RM accuracy, which measures whether the verifier accurately classifies incorrect and correct solutions. While these two metrics are correlated, RM accuracy only evaluates the verifier’s point-wise accuracy, while Best-of-N evaluates the verifier’s ability to rank solutions for choosing the correct one.
Models & Training. For training verifiers, we use open-weights Gemma models (Gemma Team et al., 2024a, b), specifically Gemma-2B for algorithmic tasks, and Gemma 2B, 7B, and Gemma-2 9B for GSM8K. For solution generation as well as LLM-as-a-Judge, we use Gemma 2B for algorithmic tasks and Gemini 1.0 Pro (Google et al., 2023) for GSM8K. For verification CoT rationales, we generate oracle rationales for algorithmic tasks programmatically (Table A.1); for GSM8K, we generate synthetic rationales using Gemini 1.0 Pro with reference-guided grading (Table A.2). See Appendix B for hyperparameter details.
4.1 Generative Verifiers Outperform Standard Verification Approaches
GenRM outperforms LLM-as-a-Judge and DPO verifiers (Figure 1), while performing comparably or slightly better than discriminative verifiers (Figure D.1). GenRM-CoT substantially improves the Best-of-N performance over GenRM. In particular, on the algorithmic tasks with oracle verification CoTs, GenRM-CoT nearly matches the oracle verifier performance.
Figure 5: Weighted Self-Consistency on MATH.
On GSM8K, GenRM-CoT consistently outperforms other methods (Figure 7, middle), even though the synthetic CoT rationales for training may contain errors. Qualitatively, GenRM-CoT is able to detect subtle reasoning errors that are missed by discriminative or direct GenRM verifiers (see Figure 2, 4, and 14).
Figure 6: Sample-Efficient Scaling with Generative Verifiers. GenRM-CoT outperforms other methods, especially for length generalization on algorithmic tasks (Gemma-2B verifiers) and easy-to-hard generalization on MATH (Gemma2-9B verifiers). Specifically, GenRM-CoT nearly matches the oracle verifier’s Best-of-N performance on algorithmic tasks. On MATH, it matches discriminative verifier’s Best-of-32 performance using 6.4×6.4$\times$6.4 × fewer solutions.
Figure 7: Easy-to-Hard Generalization on MATH, with Gemma2-9B verifiers trained only on significantly easier grade-school math problems. Compared to discriminative RMs, GenRM-CoT performs especially well on Pre-Algebra, Algebra, and Pre-Calculus, and obtains superior performance across all difficulty levels.
Easy-to-Hard Generalization. Without any training on MATH, GenRM-CoT results in a 6.4×6.4$\times$6.4 × better sample efficiency than discriminative verifiers as we increase the number of solutions to verify, and surpasses the strong self-consistency baseline (Figure 7, right). While Sun et al. (2024) demonstrate that discriminative verifiers trained on easy MATH problems can generalize to harder MATH problems, GenRM-CoT exhibits a much stronger generalization from grade-school math problems to high-school competition problems in MATH (see Figure 7 for a score breakdown by subject areas and difficulty levels).
Leveraging Self-Consistency with Verifiers. Self-consistency and test-time verification can be easily combined to boost Best-of-N performance. To do so, we use weighted self-consistency or majority-voting (Uesato et al., 2022; Liu et al., 2023; Sun et al., 2024) where we weight each solution according to the verifier’s score, and select the final answer with the largest weight (see Appendix C for details). Figure 5 shows that weighted SC can indeed improve the vanilla self-consistency (SC); in particular, weighted SC based on GenRM-CoT requires 2.5x fewer solutions than its counterpart based on Discriminative RM to reach the same performance.
4.2 Synergy Between Generation and Verification
Figure 8: SFT on correct solutions enhances verification, both for GenRM and GenRM-CoT, across all tasks. ‘Verification Only’ corresponds to verifiers trained only on verification data, by setting $\lambda=0$ in (5).
Figure 9: Unifying generation and verification boosts generation performance compared to SFT on correct solutions, in terms of Best-of-N with oracle verifier. The improvement is larger on algorithmic tasks, which use ground-truth verification data, than on GSM8K that relies on synthetic rationales, which may be inaccurate.
Unifying solution generation with verification, as done by GenRM using next-token prediction, consistently improves verification performance across all tasks, as illustrated in Figure 9. This improvement is observed for both direct and CoT-based generative verifiers, suggesting that teaching the verifier to imitate correct solutions generally helps. However, adding too much solution generation data can decrease verification performance of GenRM (Figure D.3).
Incorporating CoT verification data into the generator’s training mix leads to better solution generation performance for the GenRM-CoT verifier itself, as evidenced in Figure 9 by the improved Best-of-N scores with the oracle verifier (Pass@N). This suggests that teaching a generator to perform CoT verification using next-token prediction can deepen its understanding of the generation process itself. Overall, unifying solution generation and verification is mutually beneficial.
4.3 Scaling Model Size and Inference-time Compute
Figure 10: Scaling Inference-time Compute for Verification on GSM8K. By posing reward modeling as next-token prediction, GenRM-CoT can utilize Chain-of-Thought and Majority Voting, to turn additional test-time compute into higher percentage of problems solved under Best-of-N. Here, the horizontal line corresponds to performance of GenRM-CoT verifier with greedy decoding in Eq (6).
Figure 11: Model Scaling for Generative Verifiers. We evaluate MATH performance of Gemma 2B, 7B, and Gemma2 9B verifiers trained on GSM8K. We observe positive scaling trends for GenRM (direct) and GenRM-CoT as well as Discriminative RM, both for (Left) Best-of-N performance, and (Right) RM accuracy on the test set. Generative verifiers outperform discriminative counterparts in all model regimes.
Scaling Test-Time Compute with GenRM-CoT can be done by sampling multiple CoTs and applying majority voting, as described in Eq (7). As shown in Figure 11, GenRM-CoT verifier’s performance scales gracefully with number of votes at test time, under all three Gemma model sizes (2B, 7B, 9B), outperforming greedy decoding performance within 2 votes. Notably, across model scales, the finetuned GenRM-CoT verifier outperforms LLM-as-a-Judge , which also utilizes the same CoT approach and number of majority votes, but prompts a more capable Gemini 1.0 Pro model than Gemma models which we finetune as verifiers.
Scaling model size. In Figure 11, we show that generative verifiers, especially GenRM-CoT, exhibit better scaling behavior than discriminative RMs, both in terms of reward modeling accuracy and Best-of-N performance. Intuitively, bigger models are more capable of text generation, allowing GenRM-CoT finetuning to better tap into its chain-of-thought reasoning ability for verification. Furthermore, these results demonstrate that larger models generalize better using the same data, which matches what we expect from scaling model parameter counts under the next-token prediction loss.
Figure 12: Quality of synthetic rationales matters. Using reference guidance for synthetic rationale generation is crucial for GenRM-CoT to perform well on GSM8K: 91.7% with guidance vs. 87.8% without for Gemma-7B verifiers.
Figure 13: Quantity of synthetic rationales matter. Scaling the number of rationales per solution for GenRM-CoT on GSM8K improves both RM accuracy and Best-of-N performance. Here, we use fine-tuned Gemma-7B verifier, with greedy decoding at inference (6).
4.4 Synthetic Rationales: Quantity and Quality Matter
Our results on math reasoning tasks indicate that CoT verifiers can outperform discriminative and direct verifiers without requiring human-written verification rationales, highlighting the potential of LLM-generated rationales. We find that both the quality and quantity of these synthetic rationales matter. As shown in Figure 13, using reference-guided grading during rationale generation (§3.3) significantly improves verification performance. Furthermore, using multiple rationales per solution also improves performance, as shown in Figure 13. We suspect that this is because model-generated rationales may contain errors, such that training on multiple rationales per solution can result in an “ensembling” effect that prevents overfitting to such errors (Zhang et al., 2024).
Importantly, unlike prior work, our results on math reasoning tasks do not require a more capable model (Ankner et al., 2024; Ye et al., 2024) or humans (McAleese et al., 2024; Saunders et al., 2022) for generating verification rationales: we use the same model (Gemini 1.0 Pro) to both generate solutions to verify and synthetic verification rationales for training.
@@ 4.4 Synthetic Rationales: Quantity and Quality Matter
Figure 14: An example where GenRM-CoT catches a subtle mistake that the discriminative verifier is unable to catch. The candidate solution did not convert 90 minutes into 1.5 hours before dividing it by 7.5. However, the discriminative verifier was not able to detect this mistake likely because the solution does still appear to produce a valid-sounding percentage 90/7.5 = 12. Our proposed GenRM-CoT model is able to identify this mistake using step-by-step generative verification. The full verification output can be found in Table E.1.